A particle moves along a straight line OX. At a time t (in seconds) the distance x (in metres) of the particle from O is given by x = 40 + 12t- t3 How long would the particle travel before coming to res,


        (a)    40m


        (b)    56m


        (c)    16m


        (d)    24m


Grade:10

1 Answers

Jitender Pal
10 years ago

(b)
x = 40 + 12 t - t3
V =dx/dt = 12 - 3t2
    For V=0;  t= √(12/3)=2 sec
    So, after 2 seconds velocity becomes zero.
Value of x in 2 secs = 40 + 12.2-23
= 40 + 24 - 8 = 56 m